OpenStudy (anonymous):

Gina, Sam and Robby all rented movies from the same video store. They each rented some dramas, comedies and documentaries. Gina rented 11 movies total. Sam rented twice as many dramas, three times as many comedies, and twice as many documentaries as Gina. He rented 27 movies total. If Robby rented 19 movies total with the same number of dramas, twice as many comedies, and twice as many documentaries as Gina, how many movies of each type did Gina rent? 2 dramas, 6 comedies and 3 documentaries 1 dramas, 4 comedies and 6 documentaries 3 dramas, 5 comedies and 3 documentaries 4 dr

OpenStudy (anonymous):

You must set up a system of equations: C+D+Dr=11 3C+2D+2Dr=27 Dr+2C+2D=19

OpenStudy (loser66):



OpenStudy (loser66):

3variables, 3equations. the system can be solved!!! Good luck

OpenStudy (anonymous):

Thanks